# references.bib Electronic origin of the reorganization energy in interfacial electron transfer

### v1

@@
references.bib
Electronic origin of the reorganization energy in interfacial electron transfer
@@
The activation free energy of electron transfer (ET) reactions is governed by a crucial parameter: the reorganization energy. In heterogeneous ET at electrified solid–liquid interfaces, it is presumed that only factors in the electrolyte phase are responsible for determining the reorganization energy. Here, we demonstrate the contribution of the electronic density of states (DOS) of the electrode to the reorganization energy. Using van der Waals assembly of two-dimensional crystals, we tune the DOS of graphene and measure its impact on outer-sphere ET. We find the ensuing variation in ET rate arises from modulation in a reorganization energy associated with image potential localization in the electrode, which is dependent on the DOS. This work establishes a fundamental role of the electrode electronic structure in interfacial charge transfer.
@@
Underpinning energy conversion and chemical transformations in both biological [Moser1992Nature, Gray1996electron, Reece2009PCET] and abiological [catalysis_report, Lewis2006PNAS, Gray2009NatChem, Koper2011JEC, schmickler] systems are electron transfer (ET) processes. The efficiency of these natural or artificial ET architectures is highly dependent on the ability of the system to accommodate a desired rate of ET within an accessible or optimal range of driving force (the difference in Gibbs free energy of reactants and products). Marcus theory [Marcus1, marcus1957theory, marcus_theory_1956] provides a foundational microscopic framework for understanding the rate of ET in terms of the reaction driving force and the reorganization energy—the energetic cost required to distort the atomic configuration and solvation environment of the reactant species to resemble those of the product state. Initially formulated for homogeneous ET, involving redox-active ions in solution, extensions in the so-called Marcus–Gerischer [Gerischer_1960, Gerischer_1961] and Marcus–Hush [schmickler, hush1] formalisms rationalized heterogeneous ET processes at electrode–electrolyte interfaces, specifically addressing the ET rate constant in the weak coupling limit. Adaptation of the Marcus–Hush model by Chidsey [chidsey1991, MHC_Oldham2011] explained the dependence of interfacial ET rates on driving force and temperature by incorporating the Fermi–Dirac distribution of occupied electronic states in the electrode. In these and later [Kurchin2020MHC, yu_tunable_2022] extensions that incorporated a non-uniform (energy-dependent) density of electronic states (DOS) profile of the electrode, the DOS of the electrode serves exclusively to dictate the number of thermally accessible channels for ET. Additionally, the reorganization energy penalty is presumed to arise largely from nuclear reconfigurations in the electrolyte phase (typically those of the solvent and, to a much lesser extent, the redox molecule itself).
@@
Enhancements in interfacial charge transfer at electrodes/photoelectrodes due to electrostatic variations in carrier doping or defects (including vacancies, step edges, and grain boundaries) are commonly explained as arising from increases in the electrode DOS near the Fermi level (), ostensibly due to increased number of thermally accessible channels for ET [defects_Seh_Science2017, defects_Jaramillo_HER_Science2007, Linsey_Science2016_OER, Cheng_NatComm2023_CO2]. Yet the Marcus–Hush–Chidsey/Marcus–Gerischer framework often fails to quantitatively predict interfacial ET rate constants, even for relatively simple electrode reactions, overestimating ET rates by an order of magnitude or more[Mlimit_Koval1990, Mlimit1, Mlimit4, Mlimit5, Mlimit6, Mlimit7, Mlimit8, Mlimit_Compton2011]. A recent example of these discrepancies is found in the interfacial ET behavior of twisted bilayer graphene and twisted trilayer graphene [yu_tunable_2022, zhang_anomalous_2023]. In these moiré electrode systems, which display periodic spatial localization of the electronic charge density in moiré superlattice topological defects, models derived from the Marcus–Hush–Chidsey (MHC) framework are unable to account for the large variation in ET rate with twist angle despite modifications to account for a DOS profile that varies with energy [Kurchin2020MHC] as well as quantum capacitance effects that lead to changes in DOS of the electrode upon electrochemical polarization [yu_tunable_2022, zhang_anomalous_2023, qc1, qc5].
@@
What is missing from these frameworks to explain how electrode defects produce such strong local enhancements in interfacial ET is a consideration of the electrode’s contribution to the reorganization energy. Yet, molecular dynamics simulations have long predicted that reorganization energies for interfacial ET would vary with the distance of a redox molecular ion from a metallic electrode, owing to image charge interactions [Mlimit1, Reed2008electrochem, willard_water_2008, limaye2020understanding]. Moreover, recent simulations of twisted bilayer graphene have established a connection between the moiré twist angle and the screening [ashcroft1976, Ltf1, Ltf2] of charge carriers within the electrode, identifying that a twist-angle dependent reorganization energy can account for the interfacial ET behavior observed at moiré graphene electrodes [CoelloEscalante_TBG].
@@
Isolating two-dimensional (2D) crystals enables the assembly of vdW heterostructures with tailored electronic and chemical properties [geim_van_2013, bediako2018, DehuiDeng_catalysis2016, Huanyu_2Dforelectrocatalysis_2018]. When 2D crystals with disparate work functions are interfaced, the resulting electric field leads to charge transfer and interfacial doping, analogous to the effect of an electrostatic gate in a field effect transistor, and prior work has shown that -RuCl3 and WSe2 can thus be used to dope graphene with holes and electrons, respectively [RuCl3_KenBurch2020, Nessralla_2023]. This approach provides a modular doping mechanism similar to vacancy or substitutional doping but avoids introducing chemical disorder into the active layer. Such heterostructures provide an exceptionally well defined platform for examining how doping-induced DOS changes impact the rate of ET.
We fabricated mesoscopic electrochemical devices comprising monolayer graphene (MLG) on RuCl3. We also fabricated samples with hexagonal boron nitride (hBN) spacers placed between MLG and RuCl3, and varied the hBN spacer thickness from 3 nm to 120 nm (see Materials and Methods for additional details on sample fabrication). Electrochemical measurements were conducted using scanning electrochemical cell microscopy (SECCM), which enables nanoscale electrochemical measurements by positioning an electrolyte-filled nanopipette over the sample and forming a confined electrochemical cell upon meniscus contact [Unwin2013seccm]. In this study, we employed quartz nanopipettes of 600–800 nm diameter (Supplementary Fig. S3) containing 2 mM hexaammineruthenium(III) chloride and 100 mM potassium chloride as supporting electrolyte.
@@
Finite-element simulations were performed using COMSOL Multiphysics v.5.4 to simulate voltammetric responses with the Butler–Volmer model and estimate standard electrochemical rate constants (), as detailed in the Supplementary Information. Figure 1d illustrates the correlation between values and hBN thickness for RuCl3 (solid-state acceptor) and WSe2 (solid-state donor), revealing a strong modulation in ET kinetics with varying hBN thickness. Even when a 50 nm thick hBN spacer layer is used, RuCl3 leads to a measurable increase in relative to the electrochemical response measured in the absence of RuCl3. In the case of WSe2, the effect diminishes beyond an hBN spacer thickness of 20 nm. In the absence of an hBN spacer, MLG/RuCl3 achieves ET rates comparable to those of pristine bulk graphite.
@@
Raman spectroscopy and Hall measurements were used to quantify doping as a function of hBN thickness [raman_Das2008, Nessralla_2023, RuCl3_KenBurch2020]. Figure 2a (top) compares the Raman G-peak positions of pristine MLG with those of MLG/RuCl3, MLG/WSe2, and MLG/hBN/RuCl3 heterostructures across hBN thicknesses (Figure 2a, bottom). The doping-induced shifts in the graphene G-peak position are stronger for RuCl3 than WSe2 and are seen to attenuate with increasing hBN thickness. Prior work has shown a linear relationship between G peak position and charge carrier density [RuCl3_KenBurch2020], establishing a doping change of about cm-2 carriers per wavenumber shift. This correlation was used to deduce the graphene charge carrier density induced by RuCl3 or WSe2 heterolayers as a function of hBN spacer thickness.
@@
Figure 2c plots in MLG as a function of hBN thickness, derived from Raman G-peak shifts (circles) and Hall measurements (squares). These experimental values are compared to the doping expected from first-principles calculations [Bokdam2013] (black dashed line) that consider the work function difference between MLG and RuCl3 and dielectric constants of the spacers. While the model aligns very well with experimental data for hBN thicknesses 20 nm, doping levels exceed those predicted by theory for thinner hBN spacer layers. The enhanced doping observed in ultrathin hBN may be a result of valence band alignment between hBN and RuCl3, which activates a theoretically proposed defect-mediated charge transfer process [fludefects]. To investigate this possibility of hBN defect-mediated doping, we conducted fluorescence measurements on hBN-RuCl3 heterostructures (Figure 2e). Prior work has shown that organic solvents enhance hBN defect fluorescence via solvent-defect charge transfer [fludefects2, fludefects3]. Using wide-field imaging (see Methods), we observed that RuCl3 strongly quenches the defect-based fluorescence in hBN (Figure 2f). This result suggests that defect-mediated interactions may indeed contribute to the increased extent of MLG doping by RuCl3, which is observed when hBN 20 nm thick is used in MLG/hBN/RuCl3 heterostructures.
@@
By relating the hBN-modulated carrier densities (Figure 2c) to shifts in the chemical potential, , of MLG, Figure 3a compares experimental outer-sphere standard ET rate constants (symbols) as a function of to theoretical models (lines). In our modeling (see Materials and Methods), rate constants are normalized to those at the CNP, and are calculated at zero overpotential. Under these conditions, quantum capacitance effects on are negligible compared to the effects of solid-state doping.
@@
First, we consider an MHC model (red dotted line in Figure 3a) in which the reorganization energy (metal) is assumed to be a constant = 0.82 eV, as measured for this redox pair on metal (gold) electrodes [smalley2003heterogeneous]. This MHC model accounts for the doping dependent DOS but captures only the effect on interfacial ET of increasing the number of thermally accessible states in the electrode. The result is a prediction of very modest rate enhancement with DOS relative to that at the CNP ( 5–10-fold at extreme doping) that completely fails to replicate the experimental scaling.
@@
Instead, we find that the strongest contribution to the ET rate enhancement with increasing DOS is derived from an attenuation in the reorganization energy. This DOS dependence of arises in response to changes in the dielectric screening within the electrode as a function of doping, and can be understood in terms of image charge interactions at the electrochemical interface. To explain this effect, Figure 3b presents a series of simulation snapshots of the instantaneous discharging of an empty capacitor, displaying logarithmic maps of the induced charge distributions in response to a monovalent ion placed 5 Å away from the surface. The simulations were performed subject to a constant potential constraint, where electron-electron screening was modeled within Thomas-Fermi theory [thompson_lammps_2022, scalfi_semiclassical_2020, ahrens-iwers_electrode_2022]. The screening length, , is inversely proportional to the Fermi level, , reflecting enhanced metallicity as the charge carrier density increases. As increases (corresponding to an increased metallicity and decrease in ), the induced charge density becomes sharply localized, whereas a decrease in doping, or equivalently a decrease in metallicity, leads to a more diffuse charge distribution.
@@
Next, in Figure 3c, we account for this to model a reorganization energy, , that is a function of DOS and using non-local dielectric continuum theory. In our model, a strong attenuation of is predicted with increasing metallicity (increasing DOS), culminating in a value that converges to , consistent with metal [smalley2003heterogeneous]. This trend arises because our model predicts enhanced stabilization of the transition state during charge transfer due to sharply localized polarization in the electrode. This manifests in the calculated free energy surfaces for the Ru3+/Ru2+ redox system at different values of (left inset of Figure 3c). These calculations reveal a pronounced influence of electrode screening on electron transfer kinetics: reduced monotonically increases the activation free energy (), corresponding to the intersection of the reactant and product free energy surfaces. The right inset of Figure 3c plots vs. . The plot reveals a rapid decline in as the system is doped away from the CNP (), directly linking doping-induced Fermi level shifts to screening efficiency. Now with our model that includes —which decreases with DOS as metallicity rises (Figure 3c)—we observe quantitative agreement with experiment (red line in Figure 3a), establishing that electrode electronic structure governs and by extension the activation free energy landscape. Therefore, the effect of decreasing is the predominant contribution to increasing ET rates with increasing DOS, not the change in number of thermally accessible electron donor/acceptor states.
@@ Conclusions
These results from experimental ET measurements and theoretical calculations demonstrate the pivotal influence of the electrode electronic structure on interfacial ET kinetics and, specifically, the reorganization energy. By systematically modulating the DOS in MLG through crystalline dopants and hBN-thickness-controlled dielectric screening, we observed ET rate constant variations that deviate starkly from MHC predictions. Our findings challenge conventional theoretical frameworks, which attribute ET rate modulation by electrode DOS solely to the number of thermally accessible states at the electrode. Instead, we establish that the reorganization energy—particularly at low charge carrier densities characteristic of low-dimensional materials and semiconductors (both bulk and nanocrystalline)—is strongly governed by the electronic structure of the electrode, scaling inversely with DOS.
This paradigm shift challenges the traditional electrolyte-centric perspective of reorganization dynamics, compelling a revised activation free energy framework that explicitly integrates the influence of electrode electronic properties on . We expect this concept will also be especially relevant to photo-induced ET processes, which necessarily involve semiconducting (thus low DOS) materials [Caruge_Bawendi2008, Kamat2012, Zhu2016]. Our findings not only advance the mechanistic understanding of heterogeneous ET but also provide a foundational strategy for engineering interfacial charge transfer in next-generation devices. Future theoretical and experimental efforts should further unravel the interplay between electrode DOS, defect engineering, and to fully harness the potential of low-DOS materials in quantum technologies and energy applications. More generally, this work motivates a broader exploration of how local electronic polarizability impacts chemical reactivity (in both biological and artificial systems) through stabilization of charged transition states.

### v2

# Electronic origin of reorganization energy in interfacial electron transfer

@@
Electronic origin of reorganization energy in interfacial electron transfer
@@
Electron transfer (ET) reactions underpin energy conversion and chemical transformations in both biological[1, 2, 3] and abiological[4, 5, 6, 7] systems. The efficiency of any ET process relies on achieving a desired ET rate within an optimal driving force range. Marcus theory[8, 9] provides a microscopic framework for understanding the activation free energy, and thus the rate, of ET in terms of a key parameter: the reorganization energy. For electrified solid–liquid interfaces, it has long been conventionally understood that only factors in the electrolyte phase are responsible for determining the reorganization energy and the electronic density of states (DOS) of the electrode serves only to dictate the number of thermally accessible channels for ET.[10, 7, 11, 12, 13, 14] Here we show instead that the electrode DOS plays a central role in governing the reorganization energy, far outweighing its conventionally assumed role. Using atomically layered heterostructures, we tune the DOS of graphene and measure outer-sphere ET kinetics. We find the ensuing variation in ET rate arises from strong modulation in a reorganization energy associated with image potential localization in the electrode. This work redefines the traditional paradigm of heterogeneous ET kinetics, revealing a deeper role of the electrode electronic structure in interfacial reactivity.
@@
In the initial formulation of Marcus theory for homogeneous ET involving redox-active ions in solution, the activation free energy was explained in terms of a reorganization energy () penalty required to distort the atomic configuration and solvation environment of the reactant species to resemble those of the product state[8, 9]. Extensions in the so-called Marcus–Gerischer[10] and Marcus–Hush[7, 11] formalisms rationalized heterogeneous ET processes at electrode–electrolyte interfaces, specifically addressing the ET rate constant in the weak coupling limit. The quantum mechanical theory of ET, pioneered by Levich, Dogonadze, Chernenko, and Kuznetsov [13, 14], similarly leads to Marcus–Hush-type rate expressions. The seminal adaptation of the Marcus–Hush model by Chidsey[12] explained the dependence of interfacial ET rates on driving force and temperature by incorporating the Fermi–Dirac distribution of occupied electronic states in the electrode. In all these and later[15, 16] extensions that incorporated a non-uniform (energy-dependent) density of electronic states (DOS) profile of the electrode, the DOS of the electrode serves exclusively to dictate the number of thermally accessible channels for ET. Additionally, consistent with the original framework, is presumed to arise largely from nuclear reconfigurations in the electrolyte phase (typically those of the solvent and in some cases the redox molecule itself).
@@
Enhancements in interfacial charge transfer at electrodes/photoelectrodes due to electrostatic variations in carrier doping or defects (including vacancies, step edges, and grain boundaries) are commonly explained as arising from increases in the electrode DOS near the Fermi level (), ostensibly due to increased number of thermally accessible channels for ET[17, 18, 19]. Yet the Marcus–Hush–Chidsey/Marcus–Gerischer framework often fails to quantitatively predict interfacial ET rate constants, even for relatively simple electrode reactions, overestimating ET rates by an order of magnitude or more[20, 21, 22, 23]. A recent example of these discrepancies is found in the interfacial ET behavior of twisted bilayer graphene and twisted trilayer graphene[16, 24]. In these moiré electrode systems, which display periodic spatial localization of the electronic charge density in moiré superlattice topological defects, models derived from the Marcus–Hush–Chidsey (MHC) framework are unable to account for the large variation in ET rate with twist angle despite modifications to account for a DOS profile that varies with energy[15] as well as quantum capacitance effects that lead to changes in DOS of the electrode upon electrochemical polarization[16, 24].
@@
What is missing from these frameworks to explain how electrode defects produce such strong local enhancements in interfacial ET is a consideration of the electrode’s contribution to . Yet, molecular dynamics simulations have long predicted that for interfacial ET would vary with the distance of a redox molecular ion from a metallic electrode, owing to image charge interactions[21, 25, 26]. First-principles calculations of ET rates at graphite electrodes also considered a contribution of the electrode dielectric response to , but the effect was presumed to be sufficiently small to be neglected[27]. Moreover, recent simulations of twisted bilayer graphene have established a connection between the moiré twist angle and the screening of charge carriers within the electrode, identifying that a twist-angle dependent reorganization energy can account for the interfacial ET behavior observed at moiré graphene electrodes[28, 29]. A unique aspect of these electrode systems is the ability to continuously change the density of states at the Fermi energy, and correspondingly tune the ability of the electrode to screen electric fields. The Thomas-Fermi screening length, , quantifies the lengthscale over which charges are screened in imperfect metals. Because scales inversely with DOS, higher metallicity leads to sharper charge localization, whereas lower metallicity yields a more diffuse charge distribution. Such tunability offers a new avenue to investigate how electronic screening shapes interfacial ET.[30]
@@
Isolating two-dimensional (2D) crystals enables the assembly of vdW heterostructures with tailored electronic and chemical properties[31, 32, 33, 34]. When 2D crystals with disparate work functions are interfaced, the resulting electric field leads to charge transfer and interfacial doping, analogous to the effect of an electrostatic gate in a field effect transistor, and prior work has shown that -RuCl3 and WSe2 can thus be used to dope graphene with holes and electrons, respectively[35]. This approach provides a modular doping mechanism similar to vacancy or substitutional doping but avoids introducing chemical disorder into the active layer. Such heterostructures provide an exceptionally well-defined platform for examining how doping-induced DOS changes impact the rate of ET.
We fabricated mesoscopic electrochemical devices comprising monolayer graphene (MLG) on RuCl3. We also fabricated samples with hexagonal boron nitride (hBN) spacers placed between MLG and RuCl3, and varied the hBN spacer thickness from 3 nm to 120 nm (see Materials and Methods for additional details on sample fabrication). Electrochemical measurements were conducted using scanning electrochemical cell microscopy (SECCM), which enables nanoscale electrochemical measurements by positioning an electrolyte-filled nanopipette over the sample and forming a confined electrochemical cell upon meniscus contact[36]. In this study, we employed quartz nanopipettes of 600–800 nm diameter (Extended Data Fig. S3) containing 2 mM hexaammineruthenium(III) chloride and 100 mM potassium chloride as supporting electrolyte.
@@
Finite-element simulations were performed using COMSOL Multiphysics v.5.4 to simulate voltammetric responses with the Butler–Volmer model and estimate standard electrochemical rate constants (), as detailed in the Methods. Figure 1d illustrates the correlation between values and hBN thickness for RuCl3 (solid-state acceptor) and WSe2 (solid-state donor), revealing a strong modulation in ET kinetics with varying hBN thickness. Even when a 50 nm thick hBN spacer layer is used, RuCl3 leads to a measurable increase in relative to the electrochemical response measured in the absence of RuCl3. In the case of WSe2, the effect diminishes beyond an hBN spacer thickness of 20 nm. In the absence of an hBN spacer, MLG/RuCl3 achieves ET rates comparable to those of pristine bulk graphite.
@@
Raman spectroscopy and Hall measurements were used to quantify doping as a function of hBN thickness[35]. Figure 2a (top) compares the Raman G-peak positions of pristine MLG with those of MLG/RuCl3, MLG/WSe2, and MLG/hBN/RuCl3 heterostructures across hBN thicknesses (Figure 2a, bottom). The doping-induced shifts in the graphene G-peak position are stronger for RuCl3 than WSe2 and are seen to attenuate with increasing hBN thickness. Prior work has shown a linear relationship between G peak position and charge carrier density[35], establishing a doping change of about cm-2 carriers per wavenumber shift. This correlation was used to deduce the graphene charge carrier density induced by RuCl3 or WSe2 heterolayers as a function of hBN spacer thickness.
@@
Figure 2c plots in MLG as a function of hBN thickness, derived from Raman G-peak shifts (circles) and Hall measurements (squares). These experimental values are compared to the doping expected from first-principles calculations[37] (black dashed line) that consider the work function difference between MLG and RuCl3 and dielectric constants of the spacers. While the model aligns very well with experimental data for hBN thicknesses 20 nm, doping levels exceed those predicted by theory for thinner hBN spacer layers. The enhanced doping observed in ultrathin hBN may be a result of valence band alignment between hBN and RuCl3, which activates a theoretically proposed defect-mediated charge transfer process[38]. To investigate this possibility of hBN defect-mediated doping, we conducted fluorescence measurements on hBN-RuCl3 heterostructures (Figure 2e). Prior work has shown that organic solvents enhance hBN defect fluorescence via solvent-defect charge transfer[39]. Using wide-field imaging (see Methods), we observed that RuCl3 strongly quenches the defect-based fluorescence in hBN (Figure 2f). This result suggests that defect-mediated interactions may indeed contribute to the increased extent of MLG doping by RuCl3, which is observed when hBN 20 nm thick is used in MLG/hBN/RuCl3 heterostructures.
@@
By relating the hBN-modulated carrier densities (Figure 2c) to shifts in the chemical potential, , (i.e., at equilibrium) of MLG relative to its band structure, Figure 3a compares experimental outer-sphere standard ET rate constants (symbols) as a function of to theoretical models (lines). In our modeling (see Materials and Methods), rate constants are normalized to those at the undoped graphene (), and are calculated at zero overpotential. Under these conditions, quantum capacitance effects on are negligible compared to the effects of solid-state doping.
@@
First, we consider an MHC model (red dotted line in Figure 3a) in which the reorganization energy (metal) is assumed to be a constant = 0.82 eV, as measured for this redox pair on metal (gold modified by self-assembled monolayer thiols) electrodes[40]. This MHC model accounts for the doping dependent DOS but captures only the effect on interfacial ET of increasing the number of thermally accessible states in the electrode. The result is a prediction of very modest rate enhancement with DOS relative to that at the undoped graphene ( 5–10-fold at extreme doping) that completely fails to replicate the experimental scaling.
@@
Instead, we find that the strongest contribution to the ET rate enhancement with increasing DOS is derived from an attenuation in the reorganization energy. This DOS dependence of arises in response to changes in the dielectric screening within the electrode as a function of doping, and can be understood in terms of image charge interactions at the electrochemical interface. To explain this effect, Figure 3b presents a series of simulation snapshots of the instantaneous discharging of an empty capacitor, displaying logarithmic maps of the induced charge distributions in response to a monovalent ion placed 5 Å away from the surface. The simulations were performed subject to a constant potential constraint, where electron-electron screening was modeled within Thomas-Fermi theory [41, 30, 42]. The screening length, , is inversely proportional to the Fermi level, , reflecting enhanced metallicity as the charge carrier density increases. As increases (corresponding to an increased metallicity and decrease in ), the induced charge density becomes sharply localized, whereas a decrease in doping, or equivalently a decrease in metallicity, leads to a more diffuse charge distribution.
@@
Next, in Figure 3c, we account for this to model a reorganization energy, , that is a function of DOS and using non-local dielectric continuum theory. Our model predicts a strong attenuation of with increasing metallicity (increasing DOS), converging to a value of , consistent with metal[40]. This trend arises from enhanced stabilization of the charge transfer transition state due to sharply localized polarization within the electrode. This manifests in the calculated free energy surfaces for the Ru3+/Ru2+ redox system at different values of (left inset of Figure 3c). These calculations reveal a pronounced influence of electrode screening on electron transfer kinetics: reduced monotonically increases the activation free energy (), corresponding to the intersection of the reactant and product free energy surfaces. The right inset of Figure 3c plots vs. . The plot reveals a rapid decline in as the system is doped away from the CNP (), directly linking doping-induced Fermi level shifts to screening efficiency. By explicitly including —which decreases with rising DOS—our model achieves quantitative agreement with experimental rates (red line in Figure 3a), establishing that electrode electronic structure governs and by extension the activation free energy landscape. Therefore, the predominant contributor to increasing ET rates with increasing DOS is the decrease in , not the change in number of thermally accessible electron donor/acceptor states.
@@ Conclusions
Our experimental and theoretical results demonstrate the pivotal influence of the electrode electronic structure on interfacial ET kinetics, not only through the expected increase in the pre-exponential factor as DOS increases, but also through a significant impact on the reorganization energy itself. This paradigm shift challenges the traditional electrolyte-centric perspective of reorganization dynamics, compelling a revised activation free energy framework that explicitly integrates the influence of electrode electronic properties on . Our findings provide a more comprehensive mechanistic understanding of heterogeneous electron transfer and establish foundational principles to guide the design and optimisation of interfacial charge transfer in next-generation devices. We anticipate that this concept will be particularly relevant to photo-induced ET processes, which inherently involve semiconducting (thus low DOS) materials[43, 44, 45]. Future theoretical and experimental efforts should further unravel the interplay between electrode DOS, defect engineering, and to fully harness the potential of low-DOS materials in quantum technologies and energy applications.